• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name danez

    Log entry time 16:01:10 on February 14, 2011

    Entry number 346878

    keyword=ROC Reset Connections Fixed

    The reset connection for ROCs 1 through 4 were all found to be disconnected. Using the schematics supplied by Jack, the setups in both arms were reformed and should be working again.

    The circuit for the left arm is located in a Level Translator in CAMAC crate 2 in RR2, while the circuit for the right arm is located in a Level Translator in the NIM crate in RR2.

    Please note that the Level Translator used for the right arm was taken from the NIM crate in the FPP rack in the left arm. Only a test wire to an oscilloscope was connected, so hopefully it was okay to use.

    A copy of this log entry has been emailed to: vasulk, yez, rom